OpenStudy (anonymous):

Find p(-3) and p(5) for the function please hurry. p(x)=4x^4+8x^3-2x^2+13x+10

OpenStudy (anonymous):

Possible answers are A. 51; 3,515 B. 113; 3,473 C. -37, 1,525 D. 61; 3,525

OpenStudy (anonymous):

I have three minutes till I have to turn this in :(

OpenStudy (anonymous):

wolfram to the rescue!

OpenStudy (anonymous):

two minutes :(

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

x=-3, y=61

OpenStudy (anonymous):

p(-3)=61

OpenStudy (anonymous):

p(5)=3525

OpenStudy (anonymous):

I hope wolfram was quick enough...

OpenStudy (anonymous):

It timed out but I am just going to email it to my teacher! Thank you so much, you're a lifesaver

OpenStudy (whpalmer4):

probably too late to help you, but here's a clever way to make evaluating these big polynomials easier, especially if you don't have a calculator. Instead of \[4x^4+8x^3-2x^2+13x+10\]write \[(((4x+8)x-2)x+13)x+10\]You can evaluate this with many fewer operations. The only catch is if you have a polynomial with a missing term, you need to pretend that there is a term there, with a 0 coefficient.

OpenStudy (anonymous):

good luck with that. Next time plan ahead.

OpenStudy (whpalmer4):

I'll do x = 6 both ways: \[(((4(6)+8)6-2)6+13)6+10 = (((24+8)6-2)6+13)6+10\]\[=((192-2)6+13)6+10 = (1140+13)6+10 = 6918+10=6928\]We did 4 multiplies and 4 add/subtract \[4(6)^4+8(6)^3-2(6)^2+13(6)+10 = 4*1296+8*216-2*36+13*6+10 = 6928\]We did 12 multiples and 4 add/subtract Horner's Method (as this is called) looks complicated when written out like this, but is actually quite easy with a calculator

OpenStudy (whpalmer4):

And you can do it by just reading along the coefficients as you get some practice. For example, evaluating at x = 2, I'd say 4*2 = 8+8=16*2=32-2=30*2=60+13=73*2=146+10=156. Bam!

OpenStudy (anonymous):

thanks!

OpenStudy (whpalmer4):

It's sure to impress the cute girls, too LOL

OpenStudy (anonymous):

Except I'm a girl

OpenStudy (whpalmer4):

Okay, the cute guys! :-)

OpenStudy (whpalmer4):

And aren't you impressed? I take it for granted that you're cute ;-)